OpenStudy (anonymous):

Ms. Wong is redecorating her office. She has a choice of 7 colors of paint, 4 kinds of curtains, 3 colors of carpet, and 2 styles of furniture. How many different ways are there to redecorate if she chooses two different colors of paint, one kind of curtain, one color of carpet, and one style of furniture? A. 168 ways B. 1,008 ways C. 1,176 ways D. 23 ways

OpenStudy (anonymous):

guess my last answer didn't help much

OpenStudy (cwrw238):

different combinations of paints = 7*6 / 2 = 21

OpenStudy (anonymous):

then multiply by the other choices

OpenStudy (cwrw238):

4 types of curtain, 3 colors carpet a nd 2 of furniture number of ways = 21 * 4 * 3 * 2 = 21 * 24 = 504 ???

OpenStudy (cwrw238):

i know - but i can't see any error... ill recheck

OpenStudy (cwrw238):

i'm a bit confused with the question - he uses two different paints when he decorates , right?

OpenStudy (anonymous):

she uses 7 different paints.... Ms. Wong is redecorating her office. She has a choice of 7 colors of paint, 4 kinds of curtains, 3 colors of carpet, and 2 styles of furniture. How many different ways are there to redecorate if she chooses two different colors of paint, one kind of curtain, one color of carpet, and one style of furniture?

OpenStudy (cwrw238):

- well i'm sure that 504 is correct - the number of combinations of 2 from 7 is 21 - if the order he picks the paints is important then its 42 but the order does not matter

OpenStudy (anonymous):

A. 168 ways B. 1,008 ways C. 1,176 ways D. 23 ways these are the answers

OpenStudy (cwrw238):

none are correct

OpenStudy (anonymous):

its online schooling one of them is correct

OpenStudy (cwrw238):

the only way it could be 1008 is if he chooses 2 different colors of paint and then only uses one of them to do the decorating - that would be 2 * 504 = 1008 - thats what i meant when i said the question is a bit confusing

OpenStudy (cwrw238):

- just because its online schooling doesn't mean its correct